# Reservoir-assisted quantum battery charging at finite temperatures

### v2

@@ Abstract
Quantum batteries, as highly efficient energy storage devices, have garnered significant research interest. A key challenge in their development is to maximize the extractable energy (ergotropy) when operating within a finite-temperature reservoir. To address this, we applied quantum feedback control to the charger and investigated the effects of fermionic and bosonic thermal reservoirs on the performance of quantum batteries, including stored energy, ergotropy, and charging efficiency, in an open environment. Our findings reveal that, regardless of the type of thermal reservoir, the system exhibits optimal charging parameters. In particular, in a fermionic thermal reservoir, increasing the environmental temperature enhances battery performance, enabling stable and efficient charging. In contrast, within a bosonic thermal reservoir, higher temperatures hinder energy storage and extraction, significantly reducing charging efficiency. Additionally, we explored the impact of battery size and found that, under a fermionic reservoir, increasing the battery size appropriately can further improve performance.
@@ I Introduction
The growing demand for energy resources and the increasing environmental challenges have sparked a growing interest in the investigation of revolutionary energy storage and supply devices. The researchers anticipate that by integrating experimental techniques for precise detection and manipulation at the qubit level, they can realize a new principle of the energy storage and supply device —quantum battery Alicki and Fannes (2013). Unlike traditional batteries, which convert chemical energy or other forms of energy storage into electricity, quantum batteries store energy based on the principles of quantum mechanics, typically in the form of excited states of quantum systems. At present, quantum batteries have been explored in a variety of physical systems, including atoms and molecules Konar et al. (2022); Beleño et al. (2024); Rojo-Francàs et al. (2024); Zhang et al. (2019); Alicki (2019), spin Liu et al. (2021); Peng et al. (2021); Chen et al. (2022); Salvia et al. (2023); Yang et al. ; Le et al. (2018); Ghosh et al. (2020, 2021); Ghosh and Sen (2022); Dou et al. (2022a); Grazi et al. (2024), transmon Dou and Yang (2023); Gemme et al. (2024); Yang and Dou (2024), and micromaser Shaghaghi et al. (2022, 2023); Rodríguez et al. (2023), etc. They are expected to be smaller, possess a higher charging power Binder et al. (2015); Rossini et al. (2020); Carrasco et al. (2022); Gyhm and Fischer (2024); Campaioli et al. (2017); Ferraro et al. (2018); Seah et al. (2021); Crescente et al. (2020); Gyhm et al. (2022); Dou et al. (2022b); Zakavati et al. (2021); Mayo and Roncaglia (2022); Gao et al. (2022); Konar et al. (2024); Pokhrel and Gea-Banacloche (2025), higher charging capacity Mondal and Bhattacharjee (2022); Julià-Farré et al. (2020); Lu et al. (2021); Crescente et al. (2022); Rodriguez et al. (2023); Santos et al. (2023); Guo et al. (2024); Alimuddin et al. (2020); Wang et al. ; Cavaliere et al. (2025), and offer a greater amount of extractable work Delmonte et al. (2021); Abah et al. (2022); Arjmandi et al. (2022); Lai et al. (2024); Andolina et al. (2019); Barra (2019); Barra et al. (2022); Aparajita et al. ; Elghaayda et al. compared to traditional batteries.
@@
Open quantum batteries, which are closer to practical application scenarios, have greater research value than ideal closed quantum batteries, but they face several challenges. For example, environmental-induced decoherence causes the stored energy in quantum batteries to spontaneously dissipate, leading to battery aging and reduced charging efficiency Pirmoradian and Mølmer (2019). Additionally, a finite-temperature environment will cause the battery to tend toward a passive state, thereby hindering the extraction of energy from the quantum battery Farina et al. (2019); Bhanja et al. (2024); Ali et al. (2024). In short, as a cutting-edge technology, the quantum battery has made progress in theory and experiment Quach et al. (2022); Gemme et al. (2022); Joshi and Mahesh (2022); Hu et al. (2022); de Buy Wenniger et al. (2023); Qu et al. (2023); Campaioli et al. (2024); Razzoli et al. (2025); Lu et al. (2025), but to realize its commercialization and large-scale application, many technical challenges remain to be overcome, including environmental decoherence and energy extraction efficiency.
@@
In this work, our objective was to design an open quantum battery charging scheme that ensures stable battery charging while transforming finite temperatures into a favorable charging factor. The battery charging process is assisted by a charger which operates under homodyne quantum feedback control. This real-time feedback enables precise control of the quantum system, allowing more energy to be transferred into the battery. On the basis of this approach, we investigated the impact of bosonic and fermionic reservoirs with finite temperatures on the performance of quantum batteries. Unlike bosonic thermal reservoirs, fermionic reservoirs are typically modeled as two-level systems Artacho and Falicov (1993); Zhang et al. (2012); Chen and You (2013); Nüßeler et al. (2020); Re et al. (2020); Barthel and Zhang (2022), with their components having discrete energy levels. Some studies are also exploring the potential advantages of using this type of reservoir Kamin et al. (2024); Song et al. (2024b); Li and Jia (2011); Ghosh et al. (2012); Brodbeck et al. (2016); Malz and Nunnenkamp (2018); Damas et al. (2023); Magazzù et al. (2024).
In our proposal, we consider both the charger and the quantum battery to be modeled as two-level systems, each embedded in its respective environment; the remainder of the paper is organized as follows. In Sec. II, the charging model of the quantum battery is introduced and related physical quantities characterizing the performance of the quantum battery are provided, such as energy storage, maximum extractable energy, and charging efficiency of the battery. In Sec. III, the charging process of the quantum battery in bosonic and fermionic reservoirs is discussed. The results show that there exist optimal charging parameters, and a fermionic reservoir at a finite temperature is more conducive to enhancing the battery’s performance. In Sec. IV, the single-particle quantum battery model is extended to a multi-particle quantum battery, and the impact of increasing the number of particles in the quantum battery on its performance is discussed. Finally, we give a summary of the present protocol in Sec. V.
@@ II DISSIPATIVE QUANTUM BATTERY MODEL
Figure 1: The quantum battery model: Both the charger $C$ and the quantum battery $B$ are modeled as two-level systems and exposed to their respective dissipative environments. The corresponding transition frequencies are $\omega_{C}$ and $\omega_{B}$, respectively. The direct coupling strength between the charger and the quantum battery is $g$. The charger is driven by an external field and subject to homodyne feedback control.
II.1 The model and the master equation of the system
We model the charger and the quantum battery as a two-level atom with an energy interval of $\omega_{\rm C}$ and $\omega_{\rm B}$ between the excited and ground states, respectively. Meanwhile, the charger and quantum battery are injected into their respective reservoirs, as shown in Fig. 1. Initially, the quantum battery is in the ground state $|g\rangle$. For convenience, we choose $\hbar=1$ in the whole paper, and we assume that the charger and quantum battery share the same transition frequency, that is, $\omega_{\rm C}=\omega_{\rm B}=\omega_{0}$. The total Hamiltonian of the system is $H=H_{0}+H_{\rm I}$, where
$H_{0}=\omega_{0}\sigma_{C}^{+}\sigma_{C}^{-}+\omega_{0}\sigma_{B}^{+}\sigma_{B}^{-}+\sum\limits_{k}\omega_{Ck}a^{{\dagger}}_{k}a_{k}+\sum\limits_{k}\omega_{Bk}b^{{\dagger}}_{k}b_{k},$ (1)
the first two terms represent the free Hamiltonian of the charger and the quantum battery, while the latter two terms correspond to the free Hamiltonian of the independent environments interacting with the charger and the battery, respectively. $\sigma_{j}^{+}$ and $\sigma_{j}^{-}$ $(j=C,B)$ are the raising and lowering operators for the corresponding systems. $a^{{\dagger}}_{k}$ $(a_{k})$, $b^{{\dagger}}_{k}$ $(b_{k})$, and $\omega_{jk}$ $(j=C,B)$ are the creation (annihilation) operators and the frequency of the $k$th mode of the corresponding environments. In the interaction picture, the interacting Hamiltonian involved in the system is $H_{\rm I}=H_{\rm SI}+H_{\rm EI}$, where
$H_{\rm SI}$ $=$ $g(\sigma_{C}^{+}\sigma_{B}^{-}+\sigma_{C}^{-}\sigma_{B}^{+}),$ (2)
$H_{\rm EI}$ $=$ $\sum\limits_{k}g_{Ck}(\sigma_{C}^{+}a_{k}e^{i(\omega_{0}-\omega_{k})}+\sigma_{C}^{-}a_{k}^{{\dagger}}e^{-i(\omega_{0}-\omega_{k})})$ (3)
$+\sum\limits_{k}g_{Bk}(\sigma_{B}^{+}b_{k}e^{i(\omega_{0}-\omega_{k})}+\sigma_{B}^{-}b_{k}^{{\dagger}}e^{-i(\omega_{0}-\omega_{k})}),$
where $g$ stands for the coupling strength between the charger and the quantum battery, and $g_{jk}$ $(j=C,B)$ is the coupling strength of the system with the respective $k$th mode environment. For simplicity, we assume $\omega_{Ck}=\omega_{Bk}=\omega_{k}$. During the charging process, a resonant drive is applied to the charger and then the charger resonates with the battery to exchange energy through the interaction. In the interaction picture and considering the rotating-wave approximation, the corresponding driving Hamiltonian is
@@ II.1 The model and the master equation of the system
To improve the charging efficiency of the battery, a feedback control is applied to the charger. The photons spontaneously emitted by the charger are collected and measured by a homodyne interferometer, thus obtaining the corresponding photocurrent represented by an appropriately normalized and shifted measurement record Wiseman and Milburn (1993); Wiseman (1994); Carmichael (2009).
@@
where $\langle\sigma_{C}^{x}\rangle$ stands for the corresponding expected value. $dw(t)$ represents the Wiener increment, which satisfies $[dw(t)]^{2}=dt$ and $E[dw(t)]=0$. $\Gamma_{C}$ is the spontaneous emission rate of the charger. $\eta$ is the total measurement efficiency, which incorporates both the fraction of collected photons $\eta_{c}$ and the detector efficiency $\eta_{d}$, so that $\eta=\eta_{c}\eta_{d}$. The feedback is achieved by applying a driving field that depends on the measured results. We consider the simplest case of direct feedback, where the drive strength is proportional to the measurement record, i.e.,
@@
where $\Omega_{0}$ is a constant drive, which is set to zero here. $f$ stands for the feedback strength, and $\tau$ expresses a small time delay in the feedback loop. Here, we consider the measurement efficiency to be $\eta=1$ and assume that the decay channel corresponding to photons collected by the detector remains at effectively zero temperature. Simultaneously, the quantum battery is considered to be in contact with a thermal reservoir at a finite temperature. In the Markovian limit of $\tau\rightarrow 0$ and weak coupling conditions Hofer et al. (2017); Potts et al. (2021); Blasi et al. (2024), the evolution of the system based on feedback control can be determined by the following master equation Wiseman and Milburn (2009)
@@
The first term of Eq. (7) describes the coherent evolution introduced by the interaction between the charger and the quantum battery. The last two items in the first line describe the coherent evolution introduced by the feedback operation to the charger, and the measurement noise fed back to the system by the driver. The two terms in the second line describe the charger’s spontaneous emission and the quantum battery’s dissipation process, respectively.
The specific forms of $\mathcal{L}_{C}\rho$ and $\mathcal{L}_{B}\rho$ are as follows: $\mathcal{L}_{C}\rho=\Gamma_{C}[\sigma_{C}^{-}\rho\sigma_{C}^{+}-1/2(\sigma_{C}^{+}\sigma_{C}^{-}\rho+\rho\sigma_{C}^{+}\sigma_{C}^{-})]$ and $\mathcal{L}_{B}\rho=\gamma^{\downarrow}_{B}[\sigma_{B}^{-}\rho\sigma_{B}^{+}-1/2(\sigma_{B}^{+}\sigma_{B}^{-}\rho+\rho\sigma_{B}^{+}\sigma_{B}^{-})]+\gamma^{\uparrow}_{B}[\sigma_{B}^{+}\rho\sigma_{B}^{-}-1/2(\sigma_{B}^{-}\sigma_{B}^{+}\rho+\rho\sigma_{B}^{-}\sigma_{B}^{+})]$. Here, we consider the case where the dynamics of the quantum battery is controlled by bosonic or fermionic thermal reservoirs, respectively. If the quantum battery interacts with a bosonic thermal reservoir, $\gamma^{\downarrow}_{B}=\Gamma_{B}(1+n_{b})$ and $\gamma^{\uparrow}_{B}=\Gamma_{B}n_{b}$, where $\Gamma_{B}$ is the dissipation rate of the quantum battery and $n_{b}=1/(e^{1/T}-1)$ is the average excitation number of the corresponding thermal reservoir, in which $T$ is the dimensionless temperature of the thermal reservoir. If the quantum battery interacts with a fermionic thermal reservoir, $\gamma^{\downarrow}_{B}=\Gamma_{B}(1-n_{f})$, $\gamma^{\uparrow}_{B}=\Gamma_{B}n_{f}$ and $n_{f}=1/(e^{1/T}+1)$. In the case of the fermionic thermal reservoir, it is easy to see that $10<n_{f}<$ 1, where $20<n_{f}<1/$ 2 ($11/2<n_{f}<$ 1) means $T>0$ ($T<0$). Meanwhile, a negative temperature $T<0$ can exhibit stationary states with population inversion Carr (2013); Braun et al. (2013); Frenkel and Warren (2015); Abraham and Penrose (2017); Struchtrup (2018); Hama et al. (2018a); Assis et al. (2019); Hama et al. (2018b); Mendonça et al. (2020). Although the temperature of the fermionic thermal reservoir can be positive or negative, it is worth noting that the bosonic reservoir cannot attain negative temperatures, resulting in $0<n_{b}<\infty$. In this work, we limit the temperature of the fermionic thermal reservoir to the positive temperature range ($T>0$) to avoid population inversion induced by the thermal reservoir effects and then compare the effects of the fermionic thermal reservoir and the bosonic thermal reservoir on battery performance.
II.2 The related performance parameter of the quantum battery

### v3

@@ Abstract
Quantum batteries, as highly efficient energy storage devices, have garnered significant research interest. A key challenge in their development is to maximize the extractable energy (ergotropy) when operating within a finite-temperature reservoir. To address this, we apply quantum feedback control to the charger and investigate the effects of fermionic and bosonic thermal reservoirs on the performance of quantum batteries, including stored energy, ergotropy, and charging efficiency, in an open environment. Our findings reveal that, regardless of the type of thermal reservoir, the system exhibits optimal charging parameters. In particular, in a fermionic thermal reservoir, increasing the environmental temperature enhances battery performance, enabling stable and efficient charging. In contrast, within a bosonic thermal reservoir, higher temperatures hinder energy storage and extraction, significantly reducing charging efficiency. Additionally, we explore the impact of battery size and find that, under a fermionic reservoir, increasing the battery size appropriately can further improve performance.
@@ I Introduction
The growing demand for energy resources and the increasing environmental challenges have sparked a growing interest in the investigation of revolutionary energy storage and supply devices. Researchers anticipate that by integrating experimental techniques for precise detection and manipulation at the qubit level, they can realize a new principle of the energy storage and supply device: the quantum battery Alicki and Fannes (2013). Unlike traditional batteries, which convert chemical energy or other forms of energy storage into electricity, quantum batteries store energy based on the principles of quantum mechanics, typically in the form of excited states of quantum systems. At present, quantum batteries have been explored in a variety of physical systems, including atoms and molecules Konar et al. (2022); Beleño et al. (2024); Rojo-Francàs et al. (2024); Zhang et al. (2019); Alicki (2019), spins Liu et al. (2021); Peng et al. (2021); Chen et al. (2022); Salvia et al. (2023); Yang et al. (2025); Le et al. (2018); Ghosh et al. (2020, 2021); Ghosh and Sen (De); Dou et al. (2022a); Grazi et al. (2024), transmon Dou and Yang (2023); Gemme et al. (2024); Yang and Dou (2024), micromaser Shaghaghi et al. (2022, 2023); Rodríguez et al. (2023), etc. They are expected to be smaller, possess a higher charging power Binder et al. (2015); Rossini et al. (2020); Carrasco et al. (2022); Gyhm and Fischer (2024); Campaioli et al. (2017); Ferraro et al. (2018); Seah et al. (2021); Crescente et al. (2020); Gyhm et al. (2022); Dou et al. (2022b); Zakavati et al. (2021); Mayo and Roncaglia (2022); Gao et al. (2022); Konar et al. (2024); Pokhrel and Gea-Banacloche (2025) and higher charging capacity Mondal and Bhattacharjee (2022); Julià-Farré et al. (2020); Lu et al. (2021); Crescente et al. (2022); Rodriguez et al. (2023); Santos et al. (2023); Guo et al. (2024); Alimuddin et al. (2020); Wang et al. (2025); Cavaliere et al. (2025), and offer a greater amount of extractable work Delmonte et al. (2021); Abah et al. (2022); Arjmandi et al. (2022); Lai et al. (2024); Andolina et al. (2019); Barra (2019); Barra et al. (2022); Aparajita et al. ; Elghaayda et al. (2025) compared to traditional batteries.
@@
Open quantum batteries, which are closer to practical application scenarios, have greater research value than ideal closed quantum batteries, but they face several challenges. For example, environmentally induced decoherence causes the stored energy in quantum batteries to spontaneously dissipate, leading to battery aging and reduced charging efficiency Pirmoradian and Mølmer (2019). Additionally, a finite-temperature environment will cause the battery to tend toward a passive state, thereby hindering the extraction of energy from the quantum battery Farina et al. (2019); Bhanja et al. (2024); Ali et al. (2024). In short, as a cutting-edge technology, the quantum battery has made progress in theory and experiment Quach et al. (2022); Gemme et al. (2022); Joshi and Mahesh (2022); Hu et al. (2022); de Buy Wenniger et al. (2023); Qu et al. (2023); Campaioli et al. (2024); Razzoli et al. (2025); Lu et al. (2025), but to realize its commercialization and large-scale application, many technical challenges remain to be overcome, including environmental decoherence and energy extraction efficiency.
@@
In this work, our objective is to design an open quantum battery charging scheme that ensures stable battery charging while transforming finite temperatures into a favorable charging factor. The battery charging process is assisted by a charger which operates under homodyne quantum feedback control. This real-time feedback enables precise control of the quantum system, allowing more energy to be transferred into the battery. On the basis of this approach, we investigate the impact of bosonic and fermionic reservoirs with finite temperatures on the performance of quantum batteries. Unlike bosonic thermal reservoirs, fermionic reservoirs are typically modeled as two-level systems Artacho and Falicov (1993); Zhang et al. (2012); Chen and You (2013); Nüßeler et al. (2020); Re et al. (2020); Barthel and Zhang (2022), with their components having discrete energy levels. Some studies are also exploring the potential advantages of using this type of reservoir Kamin et al. (2024); Song et al. (2024b); Li and Jia (2011); Ghosh et al. (2012); Brodbeck et al. (2016); Malz and Nunnenkamp (2018); Damas et al. (2023); Magazzù et al. (2024).
In our proposal, we consider both the charger and the quantum battery to be modeled as two-level systems, each embedded in its respective environment. The remainder of the paper is organized as follows. In Sec. II, the charging model of the quantum battery is introduced and related physical quantities characterizing the performance of the quantum battery are provided, such as energy storage, maximum extractable energy, and charging efficiency of the battery. In Sec. III, the charging process of the quantum battery in bosonic and fermionic reservoirs is discussed. The results show that there exist optimal charging parameters, and a fermionic reservoir at a finite temperature is more conducive to enhancing the battery’s performance. In Sec. IV, the single-particle quantum battery model is extended to a multiparticle quantum battery, and the impact of increasing the number of particles in the quantum battery on its performance is discussed. We give a summary of the present protocol in Sec. V.
@@ II DISSIPATIVE QUANTUM BATTERY MODEL
II.1 Model and the master equation of the system
We model the charger and the quantum battery as a two-level atom with an energy interval of $\omega_{C}$ and $\omega_{B}$ between the excited and ground states, respectively. Meanwhile, the charger and quantum battery are injected into their respective reservoirs, as shown in Fig. 1. Initially, the quantum battery is in the ground state $|g\rangle$. For convenience, we choose $\hbar=1$ throughout the paper, and we assume that the charger and quantum battery share the same transition frequency, that is, $\omega_{C}=\omega_{B}=\omega_{0}$. The total Hamiltonian of the system is $H=H_{0}+H_{I}$, where
$H_{0}=\omega_{0}\sigma_{C}^{+}\sigma_{C}^{-}+\omega_{0}\sigma_{B}^{+}\sigma_{B}^{-}+\sum\limits_{k}\omega_{Ck}a^{{\dagger}}_{k}a_{k}+\sum\limits_{k}\omega_{Bk}b^{{\dagger}}_{k}b_{k}.$ (1)
Figure 1: Quantum battery model. Both the charger $C$ and the quantum battery $B$ are modeled as two-level systems and exposed to their respective dissipative environments. The corresponding transition frequencies are $\omega_{C}$ and $\omega_{B}$, respectively. The direct coupling strength between the charger and the quantum battery is $g$. The charger is driven by an external field and subject to homodyne feedback control.
The first two terms represent the free Hamiltonian of the charger and the quantum battery, while the last two terms correspond to the free Hamiltonian of the independent environments interacting with the charger and the battery, respectively. In addition, $\sigma_{j}^{+}$ and $\sigma_{j}^{-}$ $(j=C,B)$ are the raising and lowering operators for the corresponding systems, respectively, and $a^{{\dagger}}_{k}$ $(a_{k})$ and $b^{{\dagger}}_{k}$ $(b_{k})$, and $\omega_{jk}$ $(j=C,B)$ are the creation (annihilation) operators and the frequency of the $k$th mode of the corresponding environments. In the interaction picture, the interacting Hamiltonian involved in the system is $H_{I}=H_{SI}+H_{EI}$, with
$H_{SI}$ $=$ $g(\sigma_{C}^{+}\sigma_{B}^{-}+\sigma_{C}^{-}\sigma_{B}^{+}),$ (2)
$H_{EI}$ $=$ $\sum\limits_{k}g_{Ck}(\sigma_{C}^{+}a_{k}e^{i(\omega_{0}-\omega_{k})t}+\sigma_{C}^{-}a_{k}^{{\dagger}}e^{-i(\omega_{0}-\omega_{k})t})$ (3)
$+\sum\limits_{k}g_{Bk}(\sigma_{B}^{+}b_{k}e^{i(\omega_{0}-\omega_{k})t}+\sigma_{B}^{-}b_{k}^{{\dagger}}e^{-i(\omega_{0}-\omega_{k})t}),$
where $g$ stands for the coupling strength between the charger and the quantum battery and $g_{jk}$ $(j=C,B)$ is the coupling strength of the system with the respective $k$th mode environment. For simplicity, we assume $\omega_{Ck}=\omega_{Bk}=\omega_{k}$. During the charging process, a resonant drive is applied to the charger and then the charger resonates with the battery to exchange energy through the interaction. In the interaction picture and considering the rotating-wave approximation, the corresponding driving Hamiltonian is
@@ II.1 Model and the master equation of the system
To improve the charging efficiency of the battery, a feedback control is applied to the charger. The photons spontaneously emitted by the charger are collected and measured by a homodyne interferometer, thus obtaining the corresponding photocurrent represented by an appropriately normalized and shifted measurement record Wiseman and Milburn (1993); Wiseman (1994); Carmichael (2009)
@@
where $\langle\sigma_{C}^{x}\rangle$ stands for the corresponding expected value; $dw(t)$ represents the Wiener increment, which satisfies $[dw(t)]^{2}=dt$ and $E[dw(t)]=0$; $\Gamma_{C}$ is the spontaneous emission rate of the charger; and $\eta$ is the total measurement efficiency, which incorporates both the fraction of collected photons $\eta_{c}$ and the detector efficiency $\eta_{d}$ so that $\eta=\eta_{c}\eta_{d}$. The feedback is achieved by applying a driving field that depends on the measured results. We consider the simplest case of direct feedback, where the drive strength is proportional to the measurement record, i.e.,
@@
where $\Omega_{0}$ is a constant drive, which is set to zero here; $f$ stands for the feedback strength; and $\tau$ expresses a small time delay in the feedback loop. Here we consider the measurement efficiency to be $\eta=1$ and assume that the decay channel corresponding to photons collected by the detector remains at effectively zero temperature. Simultaneously, the quantum battery is considered to be in contact with a thermal reservoir at a finite temperature. In the Markovian limit of $\tau\rightarrow 0$ and weak-coupling conditions Hofer et al. (2017); Potts et al. (2021); Blasi et al. (2024), the evolution of the system based on feedback control can be determined by the master equation Wiseman and Milburn (2009)
@@
The first term of Eq. (7) describes the coherent evolution introduced by the interaction between the charger and the quantum battery. The last two terms in the first line describe the coherent evolution introduced by the feedback operation to the charger, and the measurement noise fed back to the system by the driver. The two terms in the second line describe the charger’s spontaneous emission and the quantum battery’s dissipation process, respectively.
The specific forms of $\mathcal{L}_{C}\rho$ and $\mathcal{L}_{B}\rho$ are as follows: $\mathcal{L}_{C}\rho=\Gamma_{C}[\sigma_{C}^{-}\rho\sigma_{C}^{+}-1/2(\sigma_{C}^{+}\sigma_{C}^{-}\rho+\rho\sigma_{C}^{+}\sigma_{C}^{-})]$ and $\mathcal{L}_{B}\rho=\gamma^{\downarrow}_{B}[\sigma_{B}^{-}\rho\sigma_{B}^{+}-1/2(\sigma_{B}^{+}\sigma_{B}^{-}\rho+\rho\sigma_{B}^{+}\sigma_{B}^{-})]+\gamma^{\uparrow}_{B}[\sigma_{B}^{+}\rho\sigma_{B}^{-}-1/2(\sigma_{B}^{-}\sigma_{B}^{+}\rho+\rho\sigma_{B}^{-}\sigma_{B}^{+})]$. Here we consider the case where the dynamics of the quantum battery is controlled by bosonic or fermionic thermal reservoirs, respectively. If the quantum battery interacts with a bosonic thermal reservoir, $\gamma^{\downarrow}_{B}=\Gamma_{B}(1+n_{b})$ and $\gamma^{\uparrow}_{B}=\Gamma_{B}n_{b}$, where $\Gamma_{B}$ is the dissipation rate of the quantum battery and $n_{b}=1/(e^{1/T}-1)$ is the average excitation number of the corresponding thermal reservoir, in which $T$ is the dimensionless temperature of the thermal reservoir. If the quantum battery interacts with a fermionic thermal reservoir, $\gamma^{\downarrow}_{B}=\Gamma_{B}(1-n_{f})$, $\gamma^{\uparrow}_{B}=\Gamma_{B}n_{f}$, and $n_{f}=1/(e^{1/T}+1)$. In the case of the fermionic thermal reservoir, it is easy to see that $10<n_{f}<$ 1, where $20<n_{f}<1/$ 2 ($11/2<n_{f}<$ 1) means $T>0$ ($T<0$). Meanwhile, a negative temperature $T<0$ can exhibit stationary states with population inversion Carr (2013); Braun et al. (2013); Frenkel and Warren (2015); Abraham and Penrose (2017); Struchtrup (2018); Hama et al. (2018a); de Assis et al. (2019); Hama et al. (2018b); Mendonça et al. (2020). Although the temperature of the fermionic thermal reservoir can be positive or negative, it is worth noting that the bosonic reservoir cannot attain negative temperatures, resulting in $0<n_{b}<\infty$. In this work, we limit the temperature of the fermionic thermal reservoir to the positive-temperature range ($T>0$) to avoid population inversion induced by the thermal reservoir effects and then compare the effects of the fermionic thermal reservoir and the bosonic thermal reservoir on battery performance.
II.2 Related performance parameter of the quantum battery